OpenStudy (anonymous):

Find the area of the larger circle if the circumference of one of the smaller circles is 36 pi mm.

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

C = pi*d 36pi = pi*d 36 = d d = 36 So the smaller circle has a diameter of 36 mm

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

this means that the diameter of the larger circle is 36+36 = 72 and the radius of the larger circle is 72/2 = 36

OpenStudy (mathstudent55):

Since you know the circumference of the small circle, you can find the radius of the small circle. Then the radius of the large circle is twice the radius of the small circle. With the radius of the large circle, you can find its area.